# Selective Excitation of Superconducting Qubits with a Shared Control Line through Pulse Shaping

### v1

@@ Abstract
In conventional architectures of superconducting quantum computers, each qubit is connected to its own control line, leading to a commensurate increase in the number of microwave lines as the system scales. Frequency-multiplexed qubit-control addresses this problem by enabling multiple qubits to share a single microwave line. However, it can cause unwanted excitation of non-target qubits, especially when the detuning between qubits is smaller than the pulse bandwidth. Here, we propose a selective-excitation-pulse (SEP) technique that suppresses unwanted excitations by shaping a drive pulse to create null points at non-target qubit frequencies. In a proof-of-concept experiment with three fixed-frequency transmon qubits, we demonstrate that the SEP technique achieves single-qubit gate fidelities comparable to those obtained with conventional Gaussian pulses while effectively suppressing unwanted excitations in non-target qubits. These results highlight the SEP technique as a promising tool for enhancing frequency-multiplexed qubit-control.
@@
A notable challenge in scaling up superconducting quantum processors is the current wiring scheme, which imposes severe space constraints within refrigerator and generates passive heat loads that can exceed the cooling capacity when the number of qubits is high [1]. Frequency-multiplexed qubit-control reduces the number of control lines needed for qubit manipulation [2, 3, 4, 5, 6, 7, 8, 9]. In this architecture, a single control line connects the electronics at room temperature to multiple qubits in the cryogenic environment. Inside the dilution refrigerator, the microwave signal is distributed to individual qubits through a divider, as illustrated in Fig. 1(a). As schematically shown in Fig. 1(b), by allocating qubit frequencies to ensure sufficient detuning relative to achievable Rabi frequencies [7, 10, 11], it is possible to suppress unwanted excitations.
Figure 1: (a) Frequency-multiplexed qubit-control architecture. Multiple qubits can be manipulated using a single microwave line. (b) By properly allocating qubit frequencies such that the spacing is wider than the bandwidth of the drive pulse, off-resonant excitation is mitigated. (c) As the number of qubits increases, the frequency spacing decreases, potentially causing the bandwidth of the drive pulse to overlap with the transition frequencies of non-target qubits. (d) Selective excitation pulse (SEP). The frequency profile of the SEP exhibits null points in the frequency-domain amplitude at the desired frequency. Aligning these null points with the frequencies of non-target qubits effectively mitigates unwanted excitations.
@@
However, this signal distribution can induce off-resonant excitations in non-target qubits when the detuning is narrower than the pulse bandwidth [Fig. 1(c)], degrading the fidelity of quantum gate operations [12]. Typically, for superconducting qubits, pulse lengths of tens of nanoseconds are used for single-qubit gate operations, resulting in a drive pulse bandwidth in the tens of megahertz range. The bandwidth of typical qubit controllers is around a few gigahertz [8], limiting the multiplexing factor to tens of qubits per control line. At this level of multiplexing, realizing a fault-tolerant quantum processor with millions of qubits would still require approximately $\mathcal{O}(10^{5})$ control lines. This already exceeds the physical limitations of the dilution refrigerator [1], implying that the degree of multiplexing is still insufficient.
@@
In this work, we propose a selective-excitation-pulse (SEP) technique to enhance the multiplexing factor of the frequency-multiplexed qubit-control architecture. The SEP technique involves designing a drive pulse that exhibits a frequency profile with null points at the frequencies of non-target qubits [13, 12, 14, 15, 16, 17, 18]. Therefore, the SEP technique enables the selective excitation of target qubits that share a single microwave line, even when the qubit frequencies are densely allocated, as illustrated in Fig. 1(d).
@@
By applying the SEP technique, we experimentally demonstrate the selective excitation of a single qubit out of three using a shared microwave line. We implement single-qubit gates based on the SEP technique and evaluate their performance using randomized benchmarking [19]. The average gate fidelity confirms the effectiveness of the SEP technique. Integrating the SEP technique with the frequency-multiplexed qubit-control architecture enhances the ability to simultaneously control a larger number of qubits using fewer microwave lines.
@@
We now outline the general concept of the SEP technique. Suppose a system comprises $N$ qubits connected to a single microwave line via a divider, as illustrated in Figs. 1(a). In this setup, when the qubits are driven by a microwave pulse at frequency $\omega_{\textrm{d}}$, the Hamiltonian in the rotating frame of the drive frequency is given by
@@
where $\Delta_{j}=\omega_{j}-\omega_{\textrm{d}}$ represents the detuning between the drive frequency $\omega_{\textrm{d}}$ and the resonance frequency of the j-th qubit $\omega_{j}$. Here, $\hbar$ is set to 1. The operators $\hat{\sigma}_{j,x},\hat{\sigma}_{j,y}$ and $\hat{\sigma}_{j,z}$ are the Pauli operators for the j-th qubit. Additionally, $s_{x}(t)$ and $s_{y}(t)$ are the real and imaginary parts of the waveform $s(t)$, respectively. When the detuning $\Delta_{j}$ is small, a drive pulse intended for a target qubit can excite non-target qubits [Fig. 1(c)].
@@
$S(\omega)=A\left[\prod_{j\in\mathcal{Q}_{\textrm{NT}}}{(\omega-\omega_{j})}\right]\exp\left[-\frac{(\omega-\omega_{d})^{2}}{2\sigma^{2}}\right],$ (2)
@@
Figure 2: (a) Frequency profiles of the control pulses. The dashed curve shows the Gaussian profile, the solid curve represents the SEP frequency profile derived from Eq. (2), and the dot-dashed curve corresponds to the profile specified in Eq. (3). Detailed parameters used to plot the frequency profiles are provided in the main text. (b) Real part $s_{x}$ and (c) imaginary part $s_{y}$ of the waveforms obtained from the frequency profiles shown in (a).
@@
Anharmonicity (MHz) $-411$ $-433$ $-413$
Energy relaxation time ($\rm{\mu}$s) $25$ $18$ $18$
Echo dephasing time ($\rm{\mu}$s) $21$ $17$ $18$
@@
$S(\omega)=A\left[\prod_{j\in\mathcal{Q}_{\textrm{NT}}}{(\omega-\omega_{j})(\omega-2\omega_{\textrm{d}}+\omega_{j})}\right]$ (3)
×$\times\exp\left[-\frac{(\omega-\omega_{\textrm{d}})^{2}}{2\sigma^{2}}\right].$
The correction transforms the frequency profile to be symmetric about the drive frequency $\omega_{\textrm{d}}$, helping to minimize phase shifts due to AC-Stark shifts [15].
@@
Now, let us consider a concrete example with two qubits, Q0 and Q1, designated as the target qubit and the non-target qubit, respectively, with frequencies $\omega_{0}$ and $\omega_{1}$. The frequency profile of the SEP, designed to suppress excitation in Q1 is shown in Fig. 2(a). Note that the corresponding frequency profiles are plotted as a function of the detuning $\Delta$ relative to the frequency of Q0. In Fig. 2(a), the solid and the dot-dashed curves represent the frequency profiles based on Eq. (2) and Eq. (3), respectively, with parameters $\sigma/2\pi=25$ MHz and $\Delta_{1}/2\pi=30$ MHz. Here, $\Delta_{1}=\omega_{1}-\omega_{0}$ represents the frequency difference between Q0 and Q1. Additionally, the frequency profile of a Gaussian pulse is shown as the dashed curve, plotted using Eq. (2) with the non-target qubit terms set to 1, with a parameter $\sigma/2\pi=25$ MHz. Note that, for all frequency profiles shown in Fig. 2(a), the amplitude $A$ is set such that the peak of each frequency profile reaches a value of 1.
@@
The real and imaginary parts of the time-domain signal, obtained from the frequency profiles shown in Fig. 2(a), are shown in Figs. 2(b) and (c), respectively. By using these signals to drive the target qubit, excitations in non-target qubits can be suppressed [20]. For more details, refer to the Supplemental Material [21].
@@
To experimentally demonstrate our proposed technique, we controlled the selected three qubits using a shared microwave line. The microwave signal for qubit control was split into three separate control lines using a power divider, with each line directed to an individual qubit [21]. While the primary motivation of this study is to reduce the number of microwave lines inside the refrigerator by installing a divider within the dilution refrigerator to split the signal just before reaching the qubits, the current experimental setup is sufficient for a proof-of-concept demonstration.
@@
In the first experiment, we implemented an $X_{\pi/2}$ gate using the SEP technique and a Gaussian pulse with a duration of 40 ns to assess the suppression of unwanted excitations in non-target qubits. We calibrated the amplitude $A$ of the frequency profile, as given in Eq. (3), to achieve a rotation angle of $\pi/2$. To compensate for phase drift caused by AC-Stark shifts during gate operations, we introduced and fine-tuned an additional detuning parameter $\delta$ of the frequency profile [21]. The waveforms used to implement the $X_{\pi/2}$ gate and the used parameters are provided in the Supplemental Material.
Using the SEP technique and Gaussian pulses, we repeatedly applied the $X_{\pi/2}$ gate to the three qubits and measured the population of each qubit. We executed 1,000 experiments to obtain $P_{0}$, the probability of finding the qubit in its ground state. After each measurement, we waited for a period of at least $10~{}T_{1}$ to ensure that the qubit state was properly initialized.
In Fig. 3, we present the Rabi oscillations observed when applying repeated $X_{\pi/2}$ gates. As shown in Figs. 3(a), (e) and (i), the $X_{\pi/2}$ gate is properly calibrated using both the SEP technique and Gaussian pulses. However, off-resonant excitations are observed in the non-target qubits when Gaussian pulses are used, as illustrated in Figs. 3(f) and (h). In contrast, the SEP technique effectively suppresses these unwanted excitations.
Next, to evaluate the average fidelity of the single-qubit gate using the SEP technique, we performed single-qubit randomized benchmarking [19]. All single-qubit Clifford gates were composed of $X_{\pi/2}$ gates and virtual-Z gates [26]. For comparison, we also evaluated the single-qubit gate fidelity using Gaussian pulses. Figures 4(a), (e) and (i) present the results of the single-qubit randomized benchmarking. Using the single-qubit gate with the SEP technique, the average single-qubit Clifford gate fidelities were determined to be 99.87(1)%, 99.80(2)%, and 99.82(2)for target qubits Q0, Q1, and Q2, respectively. In comparison, the Gaussian pulses yielded fidelities of 99.90(2)%, 99.83(2)%, and 99.83(1)for Q0, Q1, and Q2, respectively. The fidelities achieved with the SEP technique are thus comparable to those obtained with Gaussian pulses.
In addition to the gate fidelity, we examined the excitation probability for non-target qubits to quantify unwanted excitations. We fit the populations $P_{0}$ to a decay model $Ap^{L}+B$, where $p$ is the depolarizing rate. Here, the coefficients $A$ and $B$ account for state preparation errors and the asymptotic population value as the sequence length increases.
@@
Based on the results shown in Fig. 4(f)[(h)], the excitation rate is $0.2$%($0.01$%) for the SEP technique and $10$%($7$%) for Gaussian pulses. These results indicate that using the SEP technique effectively suppresses excitations compared to Gaussian pulses.
@@ Acknowledgements.
This research was supported by JST COI-NEXT (Grant No. JPMJPF2014), JST Moonshot R&\&&D (Grant No. JPMJMS2067, Grant No. JPMJMS226A), MEXT Q-LEAP (Grant No. JPMXS0118068682), JST PRESTO (Grant No. JPMJPR23F2) and JST the establishment of university fellowships towards the creation of science technology innovation (Grant No. JPMJFS2125).
Supplemental Material for
“Selective Excitation of Superconducting Qubits
with a Shared Control Line through Pulse Shaping”
I Suppression of Unintended Excitation Using Selective Excitation Pulse
To illustrate how a drive pulse based on the SEP technique suppresses unintended excitation, we analyze the dynamics of a non-target qubit under the Hamiltonian given in Eq.(2). Here, we consider a situation with two qubits: the target qubit Q0 and the non-target qubit Q1 with respective frequencies $\omega_{0}$ and $\omega_{1}$. Although our primary analysis focuses on a two-qubit system for clarity, the SEP technique and its suppression of unintended excitations can be generalized to systems with $n$ qubits. The dynamics of the non-target qubit are governed by the Hamiltonian:
$\hat{H}_{1}=-\frac{\Delta_{1}}{2}\hat{\sigma}_{1,z}+\frac{s_{x}(t)}{2}\hat{\sigma}_{1,x}-\frac{s_{y}(t)}{2}\hat{\sigma}_{1,y}.$ (5)

### v2

@@ Abstract
In conventional architectures of superconducting quantum computers, each qubit is connected to its own control line, leading to a commensurate increase in the number of microwave lines as the system scales. Frequency-multiplexed qubit control addresses this problem by enabling multiple qubits to share a single microwave line. However, it can cause unwanted excitation of non-target qubits, especially when the detuning between qubits is smaller than the pulse bandwidth. Here, we propose a selective-excitation-pulse (SEP) technique that suppresses unwanted excitations by shaping a drive pulse to create null points at non-target qubit frequencies. In a proof-of-concept experiment with three fixed-frequency transmon qubits, we demonstrate that the SEP technique achieves single-qubit gate fidelities comparable to those obtained with conventional Gaussian pulses while effectively suppressing unwanted excitations in non-target qubits. These results highlight the SEP technique as a promising tool for enhancing frequency-multiplexed qubit control.
@@
A notable challenge in scaling up superconducting quantum processors is the wiring scheme, which imposes severe space constraints within a refrigerator and generates passive heat loads that can exceed the cooling capacity when the number of qubits is high [1]. Frequency-multiplexed qubit control reduces the number of control lines needed for qubit manipulation [2, 3, 4, 5, 6, 7, 8, 9]. In this architecture, a single control line connects the electronics at room temperature to multiple qubits in the cryogenic environment. Inside the dilution refrigerator, the microwave signal is distributed to individual qubits through a divider, as illustrated in Fig. 1(a). As schematically shown in Fig. 1(b), by allocating qubit frequencies to ensure sufficient detuning relative to achievable Rabi frequencies [7, 10, 11], it is possible to suppress unwanted excitations of non-target qubits.
@@
However, this signal distribution can induce off-resonant excitations of non-target qubits, when the detuning is narrower than the pulse bandwidth [Fig. 1(c)], degrading the fidelity of quantum gate operations [12]. Typically, for superconducting qubits, pulse lengths of tens of nanoseconds are used for single-qubit gate operations, resulting in a drive pulse bandwidth in the tens of megahertz range. The bandwidth of typical qubit controllers is around a few gigahertz [8], limiting the multiplexing factor to tens of qubits per control line. At this level of multiplexing, realizing a fault-tolerant quantum processor with millions of qubits would still require approximately $\mathcal{O}(10^{5})$ control lines. This already exceeds the physical limitations of the dilution refrigerator [1], implying that the degree of multiplexing is still insufficient.
@@
In this work, we propose a selective-excitation-pulse (SEP) technique to enhance the multiplexing factor of the frequency-multiplexed qubit-control architecture. The SEP technique involves designing a drive pulse that exhibits a frequency profile $S(\omega)$ with null points at the frequencies of non-target qubits [13, 12, 14, 15, 16, 17, 18]. Therefore, the SEP technique enables the selective excitation of target qubits that share a single microwave line, even when the qubit frequencies are densely allocated, as illustrated in Fig. 1(d).
@@
By applying the SEP technique, we experimentally demonstrate the selective excitation of a single qubit out of three using a shared microwave line. We implement single-qubit gates based on the SEP technique and evaluate their performance using randomized benchmarking [19]. The average gate fidelity confirms the effectiveness of the SEP technique. Integrating the SEP technique with the frequency-multiplexed qubit-control architecture enhances the ability to simultaneously control a larger number of qubits using fewer microwave lines. We note that, while there is certainly potential for extending the SEP technique to frequency-multiplexed qubit control under conditions of frequency separations comparable to the Rabi frequency, this presents important challenges that warrant further investigation and lie beyond the scope of the present study. In particular, the calibration overhead may increase significantly with the number of qubits sharing a control line.
Figure 1: (a) Frequency-multiplexed qubit-control architecture. Multiple qubits can be manipulated using a single microwave line. (b) Qubit-frequency spacing wider than the bandwidth of the drive pulse. The off-resonant excitation is mitigated. (c) Narrow qubit-frequency spacing. The bandwidth of the drive pulse overlaps with the transition frequencies of non-target qubits. (d) Selective excitation pulse (SEP). The frequency profile of the SEP exhibits null points in the frequency-domain amplitude at the desired frequency. Aligning these null points with the frequencies of non-target qubits effectively mitigates unwanted excitations.
@@
We now outline the general concept of the SEP technique. Suppose a system comprises $N$ qubits connected to a single microwave line via a divider, as illustrated in Fig. 1(a). In this setup, when the qubits are driven by a microwave pulse at frequency $\omega_{\textrm{d}}$, the Hamiltonian in the rotating frame of the drive frequency is given by
@@
where $\Delta_{j}=\omega_{j}-\omega_{\textrm{d}}$ represents the detuning between the drive frequency $\omega_{\textrm{d}}$ and the resonance frequency of the $j$th qubit $\omega_{j}$. Here, $\hbar$ is set to 1. The operators $\hat{\sigma}_{j,x},~\hat{\sigma}_{j,y}$ and $\hat{\sigma}_{j,z}$ are the Pauli operators for the $j$th qubit. Additionally, $s_{x}(t)$ and $s_{y}(t)$ are the real and imaginary parts of the waveform $s(t)$, respectively. When the detuning $\Delta_{j}$ is small, a drive pulse intended for a target qubit may excite non-target qubits [Fig. 1(c)].
@@
$S_{\mathrm{SEP}}(\omega)=A\left[\prod_{j\in\mathcal{Q}_{\textrm{NT}}}{(\omega-\omega_{j})}\right]\exp\left[-\frac{(\omega-\omega_{d})^{2}}{2\sigma^{2}}\right],$ (2)
@@
Figure 2: (a) Frequency profiles of the control pulses. The dashed curve shows the Gaussian profile, the solid curve represents the profile $S_{\mathrm{SEP}}(\omega)$, and the dot-dashed curve corresponds to the profile $S_{\mathrm{Exp}}(\omega)$. Detailed parameters used to plot the frequency profiles are provided in the main text. (b) Real part $s_{x}$ and (c) imaginary part $s_{y}$ of the waveforms obtained from the frequency profiles shown in (a).
@@
Anharmonicity (MHz) 411-411 433-433 413-413
Energy-relaxation time ($\rm{\mu}$s) $25$ $18$ $18$
Echo dephasing time ($\rm{\mu}$s) 2121 $17$ $18$
@@
$S_{\mathrm{Exp}}(\omega)=A\left[\prod_{j\in\mathcal{Q}_{\textrm{NT}}}{(\omega-\omega_{j})(\omega-2\omega_{\textrm{d}}+\omega_{j})}\right]$ (3)
$\times\exp\left[-\frac{(\omega-\omega_{\textrm{d}})^{2}}{2\sigma^{2}}\right].$
The correction transforms the frequency profile to be symmetric about the drive frequency $\omega_{\textrm{d}}$, helping to minimize phase shifts due to ac Stark shifts [15].
@@
Now, let us consider a concrete example with two qubits, Q0 and Q1, designated as the target qubit and the non-target qubit, respectively, with frequencies $\omega_{0}$ and $\omega_{1}$. The frequency profile of the SEP, designed to suppress excitation in Q1 is shown in Fig. 2(a). Note that the corresponding frequency profiles are plotted as a function of the detuning $\Delta$ relative to the frequency of Q0. In Fig. 2(a), the solid and the dot-dashed curves represent the frequency profiles based on Eqs. (2) and (3), respectively, with parameters $\sigma/2\pi=25$ MHz and ${\color[rgb]{0,0,0}\Delta_{10}}/2\pi=30$ MHz. Here, ${\color[rgb]{0,0,0}\Delta_{10}}=\omega_{1}-\omega_{0}$ represents the frequency difference between Q0 and Q1. Additionally, the frequency profile of a Gaussian pulse is shown as the dashed curve, plotted using Eq. (2) with the non-target qubit terms set to 1, with a parameter $\sigma/2\pi=25$ MHz. Note that, for all frequency profiles shown in Fig. 2(a), the amplitude $A$ is set such that the peak of each frequency profile reaches unity.
@@
The real and imaginary parts of the time-domain signal, obtained from the frequency profiles shown in Fig. 2(a), are shown in Figs. 2(b) and (c), respectively. By using these signals to drive the target qubit, we suppress excitations in non-target qubits [20].
The SEP technique exhibits discontinuities at the initial and final points of the waveform, as shown in Fig. 2(b). These discontinuities can induce spectral leakage and associated control errors. To suppress such effects, the smoothness of the waveform is improved by appropriately tuning the spectral width parameter $\sigma$ in Eq. (2). A detailed discussion of the origin and mitigation of these discontinuities is provided in the Supplementary Material [21].
@@
To experimentally demonstrate our proposed technique, we controlled the selected three qubits using a shared microwave line. The microwave signal for qubit control was split into three separate control lines using a power divider, with each line directed to an individual qubit [21]. While the primary motivation of this study is to reduce the number of microwave lines inside the dilution refrigerator by installing a divider to split the signal just before reaching the qubits, the current experimental setup is sufficient for a proof-of-concept demonstration.
@@
In the first experiment, we implemented an $X_{\pi/2}$ gate using the SEP technique and a Gaussian pulse with a duration of 40 ns to assess the suppression of unwanted excitations in non-target qubits. We calibrated the amplitude $A$ of the frequency profile, given in Eq. (3) to achieve a rotation angle of $\pi/2$. To compensate for phase drift caused by ac Stark shifts during gate operations, we introduced and fine-tuned an additional detuning parameter $\delta$ of the frequency profile [21]. The waveforms used to implement the $X_{\pi/2}$ gate and their parameters are provided in the Supplemental Material.
Using the SEP technique and Gaussian pulses, we repeatedly applied the $X_{\pi/2}$ gates to the three qubits and measured the population of each qubit. We executed 1,000 experiments to obtain $P_{\textrm{g}}$, the probability of finding the qubit in its ground state. After each measurement, we waited for a period of at least $10~T_{1}$ to ensure that the qubit state was properly initialized.
In Fig. 3, we present the Rabi oscillations under repeated $X_{\pi/2}$ gates. As shown in Figs. 3(a), (e) and (i), the $X_{\pi/2}$ gate is properly calibrated using both the SEP and Gaussian pulses. Off-resonant excitations are observed in the non-target qubits when Gaussian pulses are used, as illustrated in Figs. 3(f) and (h). In contrast, the SEP technique effectively suppresses these unwanted excitations.
Next, to evaluate the average fidelity of the single-qubit gate using the SEP technique, we performed single-qubit randomized benchmarking [19]. All single-qubit Clifford gates were composed of $X_{\pi/2}$ gates and virtual-Z gates [26]. For comparison, we also evaluated the single-qubit gate fidelity using Gaussian pulses. Figures 4(a), (e) and (i) present the results of the single-qubit randomized benchmarking. Using the single-qubit gate with the SEP technique, the average single-qubit Clifford gate fidelities were determined to be 99.87(1)%, 99.80(2)%, and 99.82(2)% for target qubits  Q0, Q1, and Q2, respectively. In comparison, the Gaussian pulses yielded fidelities of 99.90(2)%, 99.83(2)%, and 99.83(1)% for Q0, Q1, and Q2, respectively. The fidelities achieved with the SEP technique are thus comparable to those obtained with Gaussian pulses.
In addition to the gate fidelity, we examined the excitation probability for non-target qubits to quantify unwanted excitations. We fit the populations $P_{\textrm{g}}$ to a decay model $Ap^{L}+B$, where $p$ is the depolarizing rate. Here, the coefficients $A$ and $B$ account for state preparation errors and the asymptotic population value as the sequence length increases.
@@
Based on the results shown in Fig. 4(f)[(h)], the excitation rate is 0.20.2%(0.010.01%) for the SEP technique and 1010%(77%) for Gaussian pulses. These results indicate that using the SEP technique effectively suppresses excitations compared to Gaussian pulses. Although Gaussian pulses are used as the base pulse in this study, the SEP technique is compatible with other basis-pulse shapes, including raised-cosine pulses. Additionally, SEP may be compatible with techniques such as the derivative removal by adiabatic gate (DRAG) [20, 27]. We numerically explored this possibility and found that the combination can suppress leakage to the order of $10^{-4}$ with parameters: a pulse duration $T=16\,\mathrm{ns}$, the detuning between the target qubit and the non-target qubit $46$ MHz, and an anharmonicity $\alpha/2\pi=-300\,\mathrm{MHz}$, potentially enabling faster and more robust gate implementations.
@@ Acknowledgements.
This research was supported by JST COI-NEXT (Grant No. JPMJPF2014), JST Moonshot R&\&D (Grant No. JPMJMS2067, Grant No. JPMJMS226A), MEXT Q-LEAP (Grant No. JPMXS0118068682), JST PRESTO (Grant No. JPMJPR23F2) and JST the establishment of university fellowships towards the creation of science technology innovation (Grant No. JPMJFS2125).
Supplemental Material for “Selective Excitation of Superconducting Qubits with a Shared Control Line through Pulse Shaping”
R. Matsuda,1,2,∗ R. Ohira,3,∗ T. Sumida,3 H. Shiomi,2,3 A. Machino,1,2 S. Morisaka,2,3
K. Koike,4 T. Miyoshi,2,3,4 Y. Kurimoto,3 Y. Sugita,3 Y. Ito,3 Y. Suzuki,5 P. A. Spring,6
S. Wang,6, S. Tamate,6 Y. Tabuchi,6 Y. Nakamura,6,7 K. Ogawa,2 M. Negoro2,3
1Graduate School of Engineering Science, The University of Osaka,
1-3 Machikaneyama, Toyonaka, Osaka, 560-8531, Japan
2
Center for Quantum Information and Quantum Biology, The University of Osaka,
1-2 Machikaneyama, Toyonaka, Osaka, 560-0043, Japan
3
QuEL, Inc., Hachioji ON Building 5F, 4-7-14 Myojincho, Hachioji, Tokyo, Japan
4
e-trees.Japan, Inc., Daiwaunyu Building 2F, 2-9-2 Owadamachi, Hachioji, Tokyo 192-0045, Japan
5
NTT Computer and Data Science Laboratories, NTT Corporation, Musashino 180-8585, Japan
6
RIKEN Center for Quantum Computing, Wako, Saitama 351-0198, Japan
7
Department of Applied Physics, Graduate School of Engineering,
The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-8656, Japan
S1 Suppression of Unintended Excitation Using Selective Excitation Pulse
To illustrate how a drive pulse based on the SEP technique suppresses unintended excitation, we analyze the dynamics of a non-target qubit under the Hamiltonian given by
$\hat{H}$= $\sum_{j}{\Delta_{j}\hat{a}^{\dagger}_{j}\hat{a}_{j}}+\frac{1}{2}\sum_{j}{\left[s(t)\hat{a}^{\dagger}_{j}+s^{\ast}(t)\hat{a}_{j}\right]},$ (S1)
where $\Delta_{j}=\omega_{j}-\omega_{\textrm{d}}$ represents the detuning between the drive frequency $\omega_{\textrm{d}}$ and the resonance frequency of the $j$-th qubit $\omega_{j}$. $\hat{a}^{\dagger}_{j}\,(\hat{a}_{j})$ denotes the creation (annihilation) operator of the $j$-th qubit. $s(t)$ indicates an envelope of pulse used to drive qubits. Note that we set $\hbar=1$.